OpenStudy (anonymous):

A chain weighs 10 pounds per foot. How many ounces will 4 inches weigh?

OpenStudy (texaschic101):

First you have to convert lbs to ounces and feet to inches. 1 lb = 16 ounces....so 10 lbs is 160 ounces. 1 foot is 12 inches. now make a proportion.... 160/12 = x/4 (160 ounces to 12 inches = x ounces to 4 inches) (160)(x) = (12)(4) (cross multiply) Can you solve for x ? (x will be your answer)

OpenStudy (anonymous):

1 ft= 160oz/4 in. = 40oz which would = 2.5lbs?

OpenStudy (anonymous):

or would it be 13.3 oz?

OpenStudy (texaschic101):

160/12 = x/4 I wrote it wrong the first time.... 12x = (160)(4) 12x = 640 x = 640/12 x = 53.33 ounces so if 4 inches = 53.33, then 12 inches (1 foot) = 53.33 x 3 = 159.99 when rounded = 160 ANSWER : 53.33 ounces is in 4 inches.

OpenStudy (anonymous):

Thank you So Much, I was working on it and didn't know if I had got it right, Thank you!

OpenStudy (texaschic101):

your quite welcome....thats what we are here for :)